OpenStudy (anonymous):

medal and fan to whoever can answer: y varies inversely with x. Find the constant of the inverse variation equation if x = 3.2 when y = 16. Write your answer as an integer or decimal number.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

thats the thing... i have no idea how to do this. I tried, but i just dont understand the concept

OpenStudy (skullpatrol):

Ok, the first thing to look for are the words you don't understand. What does the word "inversely" mean in the phrase "y varies inversely with x"?

OpenStudy (anonymous):

opposite, right?

OpenStudy (skullpatrol):

In this context it means y = k/x.

OpenStudy (anonymous):

okay, but what does k stand for?

OpenStudy (skullpatrol):

Where the k is the constant of the inverse variation equation.

OpenStudy (anonymous):

okay, so then what?

OpenStudy (skullpatrol):

Look for what the question is asking for...

OpenStudy (skullpatrol):

y varies inversely with x. Find the constant of the inverse variation equation if x = 3.2 when y = 16.

OpenStudy (skullpatrol):

We know the equation has the form y = k/x and we are given: "if x = 3.2 when y =16" Do you have enough info to calculate k?

OpenStudy (skullpatrol):